### v1

# Green Finance and Carbon Emissions: A Nonlinear and Interaction Analysis Using Bayesian Additive Regression Trees

Abstract
As a core policy tool for China in addressing climate risks, green finance plays a strategically important role in shaping carbon mitigation outcomes. This study investigates the nonlinear and interaction effects of green finance on carbon emission intensity (CEI) using Chinese provincial panel data from 2000 to 2022. The Climate Physical Risk Index (CPRI) is incorporated into the analytical framework to assess its potential role in shaping carbon outcomes. We employ Bayesian Additive Regression Trees (BART) to capture complex nonlinear relationships and interaction pathways, and use SHapley Additive exPlanations values to enhance model interpretability. Results show that the Green Finance Index (GFI) has a statistically significant inverted U-shaped effect on CEI, with notable regional heterogeneity. Contrary to expectations, CPRI does not show a significant impact on carbon emissions. Further analysis reveals that in high energy consumption scenarios, stronger green finance development contributes to lower CEI. These findings highlight the potential of green finance as an effective instrument for carbon intensity reduction, especially in energy-intensive contexts, and underscore the importance of accounting for nonlinear effects and regional disparities when designing and implementing green financial policies.
Keywords: green finance, carbon emissions, climate risk, BART, machine learning
1 Introduction
With the escalation of global climate change and the increasing frequency of extreme weather events, carbon reduction has emerged as an urgent global priority. Green finance, as a vital strategic instrument in addressing climate change, seeks to channel financial resources into environmentally sustainable projects aimed at improving energy efficiency and reducing carbon emissions wang2023can ; wu2024impact . It encompasses a range of financial instruments, including green credit, green bonds, and sustainable investment funds, which effectively allocate capital to projects that mitigate the adverse impacts of climate change niyazbekova2021growth . As the world’s largest carbon emitter, China has positioned green finance as a key driver in its transition toward a low-carbon economy, particularly in pursuit of its dual carbon targets: achieving carbon peaking by 2030 and carbon neutrality by 2060 lin2023china ; du2025decarbonization . According to the report Fintech Facilitates Green Finance in China: Cases and Outlook Institute2023Fintech , China has become the world’s largest green credit market and the second-largest green bond market, with the scale of green finance continuing to grow. However, the actual effectiveness of green finance is strongly shaped by local economic structures and environmental conditions. Recent studies suggest that climate risks significantly increase market uncertainty lee2024impact and influence investor decision-making behavior lu2025climate , thereby weakening the carbon reduction effects of green finance. Nevertheless, climate risk factors are rarely incorporated into analytical frameworks assessing the impact of green finance on carbon emissions. Moreover, most existing studies rely on traditional linear econometric models with manually specified interaction terms, and few have systematically examined the potential nonlinear effects or interaction mechanisms of green finance on carbon emissions.
To address these research gaps, this study employs Bayesian Additive Regression Trees (BART), which are particularly suited to capturing nonlinear relationships and complex variable interactions. Drawing on panel data covering multiple economic and environmental indicators across 30 Chinese provinces from 2000 to 2022, the analysis incorporates climate risk into the modeling framework to systematically examine how green finance affects carbon emissions and whether climate risk has a significant impact on emission outcomes.
This research makes three key contributions to bridge the gaps in variable integration and methodological innovation: first, it incorporates climate risk into the analytical framework to assess its direct effect on carbon emissions and its potential interaction effects with green finance; second, by employing BART, it overcomes the limitations of traditional linear econometric approaches in capturing nonlinear dynamics and interaction effects. Combined with interpretive tools such as Partial Dependence Plots (PDPs) and SHapley Additive exPlanations (SHAP), the study reveals the underlying mechanisms through which green finance affects carbon emissions and provides interpretable empirical evidence; third, methodologically, this study expands the analytical toolkit at the intersection of green finance and carbon emissions research, demonstrating the feasibility and value of machine learning approaches in environmental policy evaluation. It offers technical guidance for future research and provides a robust empirical foundation for the formulation of green transition policies.
2 Literature review
2.1 Impact factors of carbon emissions
Extensive and mature research has been conducted to identify the driving forces of carbon emissions. Among the various analytical frameworks, the STIRPAT model york2003stirpat remains one of the most widely applied. Yang et al. yang2018what found a temporal shift in key emission drivers from urbanization to foreign trade. Tang et al. tang2024study further highlighted that in the metal smelting industry, population, coal consumption, and urbanization were the most sensitive factors. Xiao and Peng xiao2023decomposition , through Granger causality tests, confirmed that economic growth, urbanization, industrial structure, and energy structure significantly influence carbon emissions. To better address spatial and temporal variations, Chen et al. chen2021analysis adopted the GTWR-STIRPAT model and found that energy efficiency, trade, industrial structure, and environmental regulation exerted spatially heterogeneous and time-lagged effects. Similarly, Li & Wang li2022dynamic , using the Spatial Durbin Model and Panel Threshold Model, demonstrated that energy consumption and non-green technologies drive short-term increases in carbon emissions, whereas green innovation and industrial upgrading play a key role in long-term emission reductions. In addition, a number of researchers have employed decomposition approaches to examine the structural components of carbon emissions. Chen et al. chen2019driving identified several key driving factors, including potential carbon factor effects, energy intensity effects, economic scale effects, emission efficiency effects, and energy use efficiency effects. Liu et al. liu2019driving found that economic activity, emission efficiency, and potential carbon factors were the main contributors to emission growth, while improvements in energy intensity, pure technical efficiency, and scale efficiency helped mitigate emissions. Zhang et al. zhang2022carbon concluded that economic efficiency is the sole driver of industrial carbon emissions, while energy intensity remains the most critical factor in emission reduction. In contrast, Sun et al. sun2022characteristics found that energy structure and industrial structure have relatively minor effects on the growth rate of carbon emissions, and that energy intensity is the primary factor contributing to decoupling.
2.2 Green finance and carbon emissions
Green finance, as an important financial tool for promoting low-carbon transition, has been widely confirmed by empirical studies to have a significant suppressive effect on carbon emissions. Lin et al. lin2023impact found that green finance significantly reduces local carbon emissions and generates spillover effects in surrounding regions. Among various green financial instruments, Wang et al. wang2023can argue that the emission reduction effect of green credit is more pronounced than that of green bonds. Green finance not only contributes directly through investment in low-carbon projects but also facilitates carbon reduction indirectly by stimulating green technological innovation, optimizing industrial structure, and improving energy efficiency wang2022how ; tu2025impact ; ran2023driving .
However, the emission reduction effects of green finance are not always consistent across different contexts. To explore its underlying mechanisms more deeply, recent studies have incorporated contextual variables into their analyses. Zhao and Li zhao2024role found that in regions with high political risk or high economic complexity, the effectiveness of green finance in reducing emissions is significantly weakened, suggesting that the institutional environment plays a moderating role. Ma and Fei ma2024research further emphasized that factors such as the level of governmental environmental attention and the development of the financial sector exhibit marked heterogeneity in shaping the effectiveness of green finance in influencing carbon emissions.
Moreover, researchers have begun to explore the nonlinear nature of green finance’s impact. Sun sun2022characteristics and Yu et al. yu2024machine found that the effect of green finance on carbon emissions is not a simple linear decline. Instead, it varies significantly under different conditions of energy consumption structure and financial technology development. Similarly, Cai et al. cai2025how uncovered complex nonlinear interactions between green finance and other factors such as environmental investment, inclusive finance, and financial regulation, all of which significantly influence carbon emission intensity.
In summary, while the role of green finance in carbon emission reduction is increasingly evident, its underlying mechanisms exhibit notable multidimensionality, regional heterogeneity, and nonlinearity.
2.3 Application of machine learning algorithms in carbon emission
Existing literature has made significant progress in exploring the factors influencing carbon emissions using traditional econometric models, particularly the STIRPAT framework li2017industrial ; zhang2019identifying ; thio2022estimation . However, these approaches generally rely on assumptions of linear relationships among variables and are often constrained by researchers’ prior selection of explanatory factors. As a result, they have limitations in capturing complex nonlinear relationships and high-order interaction effects. In recent years, the development of machine learning methods has provided a novel technical pathway for modeling carbon emission mechanisms and identifying key drivers. Compared to traditional models, machine learning algorithms can automatically learn nonlinear mapping relationships from large-scale datasets and exhibit greater robustness and adaptability kim2022novel ; vazquez2022automatic . As machine learning continues to expand across various domains, its application in carbon emission analysis has also gained increasing traction. For instance, various studies have applied machine learning models, such as Random Forest, Decision Trees, and XGBoost qin2022estimation ; shan2025unveiling , to identify the key drivers of carbon emissions from large sets of candidate variables, often ranging from dozens to hundreds. These approaches underscore the advantages of machine learning over traditional linear models, particularly in managing high-dimensional, complex systems and capturing nonlinear relationships and interaction effects among variables qin2022estimation ; ahmed2022influencing ; gao2023approach . To address the “black-box" nature of these machine learning algorithms, some studies have further introduced approaches to improve the interpretability of results, such as SHAP (Shapley Additive Explanations) and ALE (Accumulated Local Effects). Gao et al. gao2023approach and Shan et al. shan2025unveiling demonstrate how these methods can reveal the marginal contributions of variables and identify nonlinear “tipping points”, thereby enhancing the interpretability of model outputs and their value for policy guidance.
In summary, the impact of green finance on carbon emissions is complex, involving multiple interacting factors cai2025how . However, existing literature largely overlooks the potential nonlinear relationship between green finance and carbon emissions guo2024dataset and has yet to systematically investigate the interaction effects between green finance and other variables. These two areas remain underexplored and thus motivate our contribution.
3 Data
This research constructs a panel dataset that covers 30 Chinese provinces over the period 2000–2022. All variables are measured at the provincial level, with each variable observed separately for each province, allowing the investigation of how provincial characteristics influence carbon emission outcomes. A more detailed description of the 136 variables is provided in the following subsections.
3.1 Variable description
3.1.1 Dependent variable
We employ carbon emission intensity (CEI) as the core indicator to measure regional carbon emission performance. CEI is defined as the amount of carbon emissions per unit of GDP, reflecting the carbon efficiency of regional economic activities.
Compared to total carbon emissions, CEI controls for differences in regional economic scale, providing a more accurate measure of green development performance across regions. This makes it more suitable for analyzing the relationship between green finance development and regional carbon emissions, as it better reflects the environmental efficiency of economic activities under varying economic contexts.
Following the method proposed by Wang et al. wang2022digital , this study adopts the natural logarithm of carbon emissions per unit of GDP as the response variable. The specific formula is as follows:
$\mathrm{CEI}_{it}=\ln\frac{(\mathrm{CO}_{2})_{it}}{\mathrm{GDP}_{it}}.$
where, for province $i$ in year $t$, CEI represents the logarithmic value of carbon emission intensity, $\mathrm{CO}_{2}$ denotes total carbon emissions, and GDP denotes gross domestic product.
3.1.2 Core independent variables
Referring to Lee and Lee LEE2022105863 , Ran et al. ran2023driving and Wu et al. wu2024impact , this study constructs a composite green finance index to measure the level of green finance development across regions. The index is based on seven secondary indicators: green bond, green fund, green credit, green insurance, green equity, green support, and green investment. These sub-indicators are first standardized to ensure cross-provincial comparability, and then aggregated using the entropy weight method to generate the final composite index. A higher index value indicates a higher level of green finance development in the corresponding region. Table 1 provides additional information regarding these secondary indicators.
Table 1: Description of the core independent variables
Primary indicators Secondary indicators Variable declaration
Green Financial Index (GFI) Green bond Total green bond issuance / all bond issuance
Green fund Market Value of Green Funds / Total Market Value of All Funds
Green credit Credit for Environmental Protection Projects / Total Provincial Credit
Green insurance Revenue from Environmental Liability Insurance / Total Insurance Premium Income
Green equity Trading Volume of Carbon Emissions, Energy Use Rights, and Pollution Discharge Rights / Total Volume of Equity Market Transactions
Green support Fiscal Expenditure on Environmental Protection / General Public Budget Expenditure
Green investment Environmental pollution control investment / GDP
3.1.3 Other independent variables
Existing literature on the drivers of carbon emissions from an energy perspective has primarily focused on dimensions such as energy structure, energy efficiency, and energy intensity chen2019driving ; liu2019driving ; zhang2022carbon ; xiong2025emission . Meanwhile, total energy consumption (TEC) has been widely recognized as a key factor contributing to the increase in carbon emissions and has become a general consensus in the academic community li2022dynamic ; yu2024machine . However, when shifting the focus to carbon emission intensity, the impact of TEC becomes more complex. zhao2024role . Therefore, this paper includes TEC as one of the explanatory variables. On the one hand, it aims to further examine the impact of TEC on CEI. On the other hand, it seeks to explore whether green finance interacts with TEC in a way that facilitates the transformation of the energy consumption structure and ultimately contributes to the reduction of carbon emission intensity.
Climate change has become a central concern globally, and managing climate risk exposure through investment strategies has emerged as a key decision-making factor for investors lu2025climate . Climate risk, particularly physical climate risk, can increase market uncertainty and reduce investment incentives, thereby inhibiting the development of green finance markets. In this context, drawing on the methodology proposed by Guo et al. guo2024impact , this study constructs a Climate Physical Risk Index (CPRI) based on region-specific meteorological observations and incorporates it as one of the explanatory variables in the model. To capture both the intensity and frequency of extreme weather events, the CPRI evaluates physical climate risk across four dimensions: extreme low temperature, extreme high temperature, extreme rainfall, and extreme drought. Based on daily meteorological data from local stations, the annual number of days exceeding each threshold is calculated for every province. To ensure comparability across provinces, each of the four sub-indices is standardized. The overall CPRI is then computed as the weighted average of these standardized sub-indices. A higher CPRI value indicates greater exposure to physical climate risk in a given region.
In addition, carbon emissions are also influenced by a range of socioeconomic and environmental factors. Building on the existing literature from section 2, this study incorporates several variables into the model to better capture the potential relationships among key predictors. These include level of economic development, industrial structure, research and development (R&D) intensity, urbanization level, population density, foreign direct investment, degree of government intervention, environmental regulation intensity. In light of empirical studies that have confirmed the significant regional heterogeneity in the impact of green finance on carbon emissions, this study further divides the provinces into eastern, central, and western regions based on the traditional regional classification method, and incorporates this classification as a categorical variable in the model.
3.2 Data sources
This study constructs a panel dataset covering 30 provinces in China from 2000 to 2022. Due to severe data limitations, the Tibet, Hong Kong, Marco and Taiwan is excluded from the analysis. A more detailed description of the data is shown in Table 2.
Table 2: Data description
Index Sign Description Data source
Carbon emission intensity CEI Total Carbon emission / GDP China Energy Statistical Yearbook; Regional Statistical Yearbook
Green finance index GFI The index system is constructed from the seven dimensions of green credit, green equity, green insurance, green support, green investment, green bond, and green fund, and the weight is given by the entropy method China Energy Statistical Yearbook; Regional Statistical Yearbook; Environmental Status Bulletin; Sectoral Yearbooks by Government Departments; CSMAR database
Total energy consumption TEC Final energy consumption + energy transformation losses + energy losses China Statistical Yearbook; China Energy Statistical Yearbook; Regional Statistical Yearbook
Economic development GDP Per-capita GDP China Statistical Yearbook; China Energy Statistical Yearbook; Regional Statistical Yearbook
Foreign direct investment level FDI Foreign direct investment / GDP China Statistical Yearbook; China Energy Statistical Yearbook; Regional Statistical Yearbook
Industrial structure industry Output value of the tertiary sector / output value of the secondary sector China Statistical Yearbook; China Energy Statistical Yearbook; Regional Statistical Yearbook
Urbanization urbanize Urban population / total population China Statistical Yearbook; China Energy Statistical Yearbook; Regional Statistical Yearbook
Population density population Total regional population / administrative area size China Statistical Yearbook; China Energy Statistical Yearbook; Regional Statistical Yearbook
R&D intensity RDI Internal R&D expenditure / regional GDP China Statistical Yearbook; China Energy Statistical Yearbook; Regional Statistical Yearbook
Government intervention level govern Fiscal expenditure / regional GDP China Statistical Yearbook; China Energy Statistical Yearbook; Regional Statistical Yearbook
Environmental regulation environment Completed investment in industrial pollution control / industrial value added China Statistical Yearbook; China Energy Statistical Yearbook; Regional Statistical Yearbook
Climate Physical Risk Index CPRI Combine the four standardized sub-indices: extreme low temperature, extreme high temperature, extreme rainfall, and extreme drought using weighted average Deep Data database
3.3 Data preprocessing
To reduce right-skewness in the distribution of variables, we applied logarithmic transformations to the dependent variable (CEI) and to selected independent variables, namely economic development (GDP), population density (population), industrial structure (industry), and environmental regulation (environment), all of which exhibited skewness statistics greater than 2. We then employed Generalized Variance Inflation Factors (GVIF) to conduct a preliminary analysis of potential multicollinearity among explanatory variables. A GVIF value close to 1 indicates little to no multicollinearity, while values equal to or greater than 5 or 10 suggest strong multicollinearity obrien2007caution . Similarly, the adjusted generalized standard error inflation factor (aGSIF) values exceeding 2.2 or 3.2 are also considered indicative of significant multicollinearity (Fannie et al., 2025).
As shown in Table 3, the GVIF for the variable representing the level of economic development is 5.55 (>>5), and the corresponding aGSIF is 2.36 (>>2.2), indicating notable multicollinearity. Therefore, we have decided to exclude this variable from the final model specification to avoid estimation bias.
Table 3: Multicollinearity results
Variables GVIF aGSIF
TEC 1.73 1.31
ln(GDP) 5.55 2.36
ln(population) 4.02 2.01
ln(industry) 2.24 1.49
ln(environment) 1.50 1.22
GFI 2.43 1.56
RDI 4.08 2.03
CPRI 1.23 1.11
FDI 1.91 1.38
govern 2.62 1.62
urbanize 3.63 1.91
Region (df=2) 5.34 1.52
4 Methodology
4.1 Model specification
Originally proposed by Chipman et al. chipman2010bart , Bayesian Additive Regression Trees (BART) is a flexible, nonparametric Bayesian ensemble method that represents the response variable $y$ as the sum of $m$ regression trees:
$y=\sum_{j=1}^{m}g(x;T_{i},M_{i})+\varepsilon,\quad\varepsilon\sim N(0,\sigma^{2}),$
where, $T_{i}$ denotes the structure of the ithi-th tree, and $M_{i}$ represents its associated terminal node parameters. Each tree is constrained to be a weak learner, capturing only a small portion of the signal. To regularize the model and prevent overfitting, informative priors are placed on both the tree structures and terminal node values. These priors encourage shallow trees and shrinkage toward zero, thereby improving generalization and stabilizing the overall model.
Model estimation is carried out using a Bayesian backfitting Markov Chain Monte Carlo (MCMC) algorithm, implemented in the form of Gibbs sampler with Metropolis-Hastings updates. In each iteration, the sampler sequentially updates each tree $(T_{i},M_{i})$, conditional on the remaining m1m-1 trees. This is achieved by computing the partial residual:
$R_{i}=y-\sum_{k\neq i}g(x;T_{k},M_{k})$
which represents the part of the response not yet explained by the other trees. The residual $R_{i}$ serves as the local target for updating $T_{i}$ and $M_{i}$. By fitting each tree to the current residuals, the algorithm gradually refines the overall model, effectively adapting to nonlinear patterns and latent interaction effects in the data.
The update of the tree structure $T_{i}$ is conducted using a Metropolis–Hastings (MH) step. New tree proposals are generated through operations such as growing or pruning nodes, modifying split rules, or swapping decision nodes. The acceptance probability is computed by integrating out the terminal node parameters $M_{i}$, which is tractable due to the use of conjugate priors. Once the new tree is accepted, $M_{i}$ is updated by drawing from its full conditional posterior distribution, which follows a normal distribution. The noise variance $\sigma^{2}$ is updated at the end of each iteration by drawing from a full conditional inverse-gamma distribution. By running the MCMC algorithm for a sufficiently large number of iterations and discarding the initial samples as burn-in, a sequence of fitted functions is obtained:
$f^{(k)}(x)=\sum_{i=1}^{m}g(x;T_{i}^{(k)},M_{i}^{(k)}).$
These samples are treated as draws from the posterior distribution $p(f(x)|y)$, forming the basis for Bayesian inference. They can be used for prediction, uncertainty quantification, partial dependence analysis, and variable selection. The model framework and sampling algorithm together enable BART to flexibly capture complex nonlinear relationships and high-order interactions, even when such interactions are not explicitly specified in the model.
The model is evaluated using Root Mean Square Error (RMSE) and pseudo-$R^{2}$. A lower RMSE indicates a better model fit. The RMSE is calculated as follows:
$RMSE=\sqrt{\frac{1}{n}\sum_{i=1}^{n}(y_{i}-\hat{y}_{i})^{2}},$
The pseudo-$R^{2}$ is calculated using the formula:
$\text{pseudo-}R^{2}=1-\frac{SSE}{SST}.$
where SSE is the sum of squared errors and SST is the total sum of squares. It indicates the proportion of variation in the dependent variable that is explained by the model. A higher pseudo-$R^{2}$ suggests better explanatory power.
4.2 Interpretability methods
Despite the strong predictive performance of the BART model in both regression and classification tasks, it is often considered a “black-box" predictor due to its lack of transparency and interpretability chipman2010bart . Kapelner and Bleich kapelner2016bartmachine incorporated visualization tools such as partial dependence plots (PDPs) into the bartMachine framework. However, while PDPs provide insight into average marginal effects, they may fall short in capturing more complex or localized feature interactions.
To complement PDPs and provide a more nuanced interpretation of model behavior, this study also employs SHapley Additive exPlanations (SHAP). These tools examine feature contributions from different theoretical perspectives, thereby enhancing the depth and reliability of the model interpretation, especially for complex interactions and heterogeneous effects.
4.2.1 Partial Dependence Plot
This study employs the Partial Dependence Plot (PDP) from the bartMachine framework to examine how a specific predictor affects the response variable on average, while controlling for other predictors. The PDP is based on the formulation introduced by Friedman friedman2001greedy :
$f_{j}(x_{j})=E_{x_{-j}}[f(x_{j},x_{-j})]:=\int f(x_{j},x_{-j})dP(x_{-j}),$
Since the true model function $f$ and the marginal distribution $dP(x_{-j})$ are unknown, this expectation is approximated by averaging over the training data:
$\hat{f}_{j}(x_{j})=\frac{1}{n}\sum_{i=1}^{n}\hat{f}(x_{j},x_{-j,i})$
where $x_{-j,i}$ represents all variables except $x_{j}$ from the ithi-th training sample, and $\hat{f}$ denotes predictions generated by the bartMachine model.
where $x_{-j,i}$ represents all variables except $x_{j}$ from the ithi-th training sample, and $\hat{f}$ denotes predictions generated by the bartMachine model.
Since BART is a Bayesian method, it provides a full posterior distribution for the predictions. As a result, we can construct credible intervals for the PDP estimates. These credible bands are derived by replacing $\hat{f}$ in the above formula with functions that compute specific quantiles from the post-burn-in MCMC samples of the predicted values $\hat{y}$. This enhances both the robustness and interpretability of the PDP results.
4.2.2 SHapley Additive exPlanations
SHAP (SHapley Additive exPlanations) is grounded in the game-theoretic concept of Shapley values, which were originally proposed for model explanation by Štrumbelj & Kononenko strumbelj2011general ; strumbelj2014explaining . However, the approach only gained widespread attention after Lundberg & Lee lundberg2017unified introduced SHAP as a unified framework for feature attribution. Their work not only rebranded the method but also extended its theoretical foundations, connecting Shapley values with other model-agnostic interpretability techniques such as LIME.
SHAP treats features as “players" in a cooperative game, where the model prediction is the “payout" to be fairly distributed among them. The contribution of each feature is defined as its average marginal contribution across all possible coalitions of features. Formally, the Shapley value for feature $j$, given a prediction function $f$ and input $x$, is defined as:
$\phi_{i}=\sum_{S\subseteq F\setminus\{i\}}\frac{|S|!\cdot(|F|-|S|-1)!}{|F|!}[f(S\cup i)-f(S)]$
where $F$ is the set of all features; $S$ is the subset of features excluding $i$; $f(S)$ is the expected prediction when only features in $S$ are known.
In practice, for models with a large number of input variables, exhaustively enumerating all feature subsets S to compute Shapley values is computationally infeasible. To address this challenge, we employed the FastSHAP R package developed by Greenwell and Greenwell greenwell2020fastshap , which provides an efficient approximation of Shapley values for any supervised learning model. Rather than calculating exact Shapley values by iterating over all possible feature coalitions, FastSHAP employs a Monte Carlo-based approximation method proposed by Strumbel & Kononenko strumbelj2014explaining . The approximated outputs, commonly referred to as SHAP values, preserve the theoretical properties of Shapley values while enabling scalable interpretation of complex black-box models.
In our analytical framework, we first use Partial Dependence Plots (PDPs) from the bart package to explore the global structure and interaction patterns of the model. Then, we apply SHAP values to conduct detailed mechanism analysis. This combined approach helps provide a more comprehensive understanding of how input features affect the predictions and enhances the transparency and credibility of BART.
5 Results
5.1 Model performance
5.1.1 Model evaluation
To ensure a realistic evaluation of the model’s predictive performance and to avoid data leakage, we used data from 2000 to 2020 as the training set and data from 2021 to 2022 as the test set. Additionally, we applied 10-fold cross-validation to evaluate the model’s generalization ability. The BART model initially trained with default hyperparameters achieved a pseudo-R² of 0.989 and an RMSE of 0.1 on the training set. The 10-fold cross-validation yielded a pseudo-R² of 0.9586 and an RMSE of 0.1886. On the test set, the model achieved a pseudo-R² of 0.8493 and an RMSE of 0.2826. These results indicate that even with default settings, BART demonstrates strong fitting ability and successfully captures most of the underlying data patterns.
However, the model’s predictive performance on unseen data was relatively weaker, and the residuals failed the normality test (p-value=0.00091<0.05), indicating potential model misspecification or non-Gaussian error distribution. To address this, we used the built-in grid search function in the bartMachine package to tune the model’s hyperparameters. The tuned model showed improved performance on the training set, with a pseudo-R² of 0.9955 and an RMSE of 0.06. Importantly, the cross-validated pseudo-R² remained stable at 0.9608, suggesting stable performance without signs of overfitting. Most notably, the pseudo-R² on the test set increased from 0.8493 to 0.8548, confirming that hyperparameter tuning enhanced the model’s predictive accuracy on new data. Moreover, the residuals from the tuned model passed the normality test (p-value = 0.269 >> 0.05), indicating a better fit to the data.
In summary, the tuned BART model demonstrates both high predictive power and robust generalization. Therefore, all subsequent analyses on the relationships between green finance, climate risk, and carbon emission intensity are based on the tuned model.
5.1.2 Model Diagnostics and Statistical Significance Testing
Figure 1 presents the residuals Q-Q plot to assess normality and the residuals vs predicted values plot to assess homoskedasticity. The Shapiro-Wilk test for residuals yielded a p-value of 0.269 >> 0.05, indicating that the residual distribution generally conforms to normality. Although a few points deviate in the tails of the Q-Q plot, the majority of points align closely along the diagonal line. Meanwhile, the residuals vs predicted values plot shows no obvious pattern or trend, suggesting that there is no significant heteroskedasticity issue.
In addition, bartMachine provides a permutation-based approach to assess the statistical significance of individual predictors or groups of variables, analogous to the partial F-test in ordinary least squares regression but without assuming linearity. We applied this method to test the overall model validity, as well as the individual significance of the GFI and CPRI.
For the overall model test, we permuted the response variable $y$, thereby breaking any association between the predictors and the outcome. This is equivalent to an omnibus F-test in linear regression. The result shows a p-value of 0, well below the conventional 0.05 threshold, indicating that the model’s predictive ability is highly statistically significant.
Also, we conducted individual significance tests for GFI and CPRI. In each case, the values of the target variable were randomly permuted within the sample, destroying its relationship with the response while keeping all other variables fixed. This allows us to test whether the variable contributes predictive power after controlling for others. The results show that GFI has a p-value of 0.0396 < 0.05, indicating a statistically significant effect on carbon emission intensity. In contrast, CPRI yields a p-value of 0.238, which exceeds the 0.05 threshold. Therefore, we fail to reject the null hypothesis and conclude that there is insufficient statistical evidence to support a significant effect of CPRI on carbon emission intensity.
Also, we conducted individual significance tests for GFI and CPRI. In each case, the values of the target variable were randomly permuted within the sample, destroying its relationship with the response while keeping all other variables fixed. This allows us to test whether the variable contributes predictive power after controlling for others. The results show that GFI has a p-value of 0.02, indicating a statistically significant effect on carbon emission intensity. In contrast, CPRI yields a p-value of 0.281, which exceeds the 0.05 threshold. Therefore, we fail to reject the null hypothesis and conclude that there is insufficient statistical evidence to support a significant effect of CPRI on carbon emission intensity.
Figure 1: Test of normality of errors using $Q$-plot and the Shapiro-Wilk test (left), residual plot to assess heteroskedasticity (right).
5.2 Feature importance and interaction terms
This section utilizes the built-in functionalities of bartMachine to examine the importance of individual predictors and their interactions in BART, Additionally, SHAP methods are employed to visualize nonlinear relationships and interactions among variables. Figures 2, 3, and 4 present the variable importance and interaction results extracted from bartMachine, while Figures 5, and 6 show visualizations based on SHAP. As shown in Figure 2, TEC appears most frequently in the model’s decision paths and has relatively narrow 95% credible intervals (CI), indicating that it contributes the most to explaining variations in CEI, with a high degree of certainty. GFI ranks fifth, and its corresponding 95% CI is also relatively narrow, suggesting that green finance exerts a significant and stable influence within the model. In contrast, CPRI ranks the lowest, with a near-zero contribution, implying limited predictive relevance. The statistical significance of each variable are further validated in Section 5.1.2.
Regarding interaction effects, Figure 3 displays the top 10 variable pairs with the strongest interactions. Among them, GFI is involved in several interactions, including GFI × TEC, GFI × population, and GFI × govern. This suggests that the impact of green finance on CEI largely depends on its interaction with other key variables. Notably, the GFI × TEC interaction ranks fifth overall and is the most prominent among GFI-related pairs. Therefore, the subsequent analysis will focus on the interaction between GFI and TEC to explore the mechanism through which green finance influences the marginal effect of energy consumption on carbon emission intensity.
Figure 2: The average inclusion proportions of explanatory variables in the tuned BART model, reflecting how frequently each variable appears in the model’s decision paths. The y-axis represents the average inclusion proportions calculated over 100 model constructions. The segments atop the bars indicate the 95% confidence intervals.
Figure 3: In the tuned BART model, this figure displays the top 10 variable pairs ranked by average interaction counts, based on 25 model constructions. The vertical axis represents the average number of interactions, and the segments atop the bars indicate 95% confidence intervals.
5.3 Nonlinear and interaction effects analysis
5.3.1 Nonlinear and regional effects of GFI
Figures 4 and 5 jointly explore the nonlinear and heterogeneous effects of GFI on CEI, using PDP from BART and SHAP visualizations. Figure 4 displays the PDP of GFI, illustrating the average marginal effect of GFI on CEI while holding all other variables constant. The results reveal a clear inverted U-shaped pattern: at low levels of GFI, the PDP curve rises, indicating that the initial development of green finance is associated with an increase in CEI. However, once GFI exceeds a certain threshold (around 0.3), the curve begins to decline, suggesting that further development of green finance helps reduce CEI. This implies that green finance has greater emission reduction potential in more advanced stages. The relatively narrow 95% confidence interval further confirms the stability and robustness of this nonlinear relationship.
The SHAP-based analysis in Figure 5 further reveals regional heterogeneity in the observed nonlinear relationship. In Eastern and Central China, the inverted U-shape is more pronounced, with most observations concentrated within the GFI range of 0.2-0.4. After GFI exceeds 0.3, SHAP values in both regions show a general downward trend. However, in Central China, most SHAP values remain positive, indicating that although GFI has reached a relatively high level, its emission reduction effect has not yet fully materialized, possibly due to the green finance sector still being in its expansion or investment phase. In contrast, in Eastern China, SHAP values decline more significantly after GFI exceeds 0.3, with some observations even turning negative. This may reflect a more mature green finance system in the region, which is beginning to effectively curb carbon intensity. In Western China, GFI levels are more widely distributed. When GFI is relatively low, SHAP values are mostly negative, suggesting that even in its early stages, green finance has already contributed to reducing CEI to some extent. As GFI increases, SHAP values gradually rise and approach positive territory, indicating that the emission reduction effect weakens during the mid-development stage. However, when GFI reaches approximately 0.3, SHAP values begin to decline again, showing signs of diminishing marginal returns or a “saturation effect.”
These findings confirm that the effectiveness of green finance in reducing carbon emissions is nonlinear and regionally differentiated, highlighting the importance of region-specific strategies in promoting green finance development.
Figure 4: PDP of GFI on CEI. The black line represents the average marginal effect of GFI on CEI when all other variables are held constant. The shaded gray area and blue lines indicate the 95% credible interval, reflecting model uncertainty. The narrowing of the credible interval suggests increasing model certainty. The x-axis shows GFI values at specified quantiles (5%, 10%, …, 95%), while the y-axis represents the average predicted CEI at each GFI level, conditional on other variables being fixed.
Figure 5: SHAP PDP of GFI on CEI, color-coded by region. Each point represents an observation, with the x-axis showing the value of GFI and the y-axis showing the corresponding SHAP value, which reflects the marginal contribution of GFI to the model prediction of CEI. SHAP values represent the marginal contribution of a given feature (GFI) to the model’s prediction of CEI. A positive SHAP value indicates that GFI increases predicted CEI (i.e., contributes positively), while a negative value suggests a suppressing effect (i.e., contributes negatively). The color of the points denotes regional classification: Eastern (red), Central (purple), and Western (yellow).
5.3.2 Interaction effects between GFI and TEC
Figure 6 presents SHAP interaction effect plots between TEC and GFI, aiming to gain insights into complex feature interactions and uncover latent patterns. The left panel shows that when TEC is relatively low (below 10000) and GFI is also at a low level, the SHAP values are predominantly positive. This indicates that under conditions of low energy consumption and limited green finance development, carbon emission intensity tends to increase. As TEC increases beyond this level, accompanied by rising GFI levels, the SHAP values exhibit a clear downward trend and eventually turn negative. This suggests that when energy consumption is high, the interaction with higher levels of green finance leads to a negative marginal effect on CEI. In other words, the expansion of green finance may help decouple energy consumption from emissions, alleviating the environmental pressure typically associated with high energy use. In the right panel, as both TEC and GFI increase, the SHAP values shift from yellow to purple. This color gradient visually reinforces the finding that the marginal harm of energy consumption on carbon intensity becomes weaker, or even beneficial under high GFI conditions.
Taken together, the interaction between GFI and TEC reveals a clear nonlinear and synergistic relationship in shaping carbon emission intensity. Green finance not only directly influences CEI but also plays a critical moderating role by mitigating the environmental cost of energy use. These findings emphasize the importance of integrating financial instruments into carbon mitigation strategies. In particular, energy-intensive regions stand to benefit more from targeted green finance policies that can curb emissions even amid rising energy demand. Moreover, accounting for such interaction effects contributes to a more nuanced understanding of how various policy instruments, including financial incentives, energy regulation and regional planning, can be effectively coordinated to achieve long-term environmental objectives.
In summary, the interaction between population and GFI in shaping carbon emission intensity is highly complex and stage-dependent. These findings highlight the need for regionally differentiated green finance policies and call for particular attention to the potential “saturation effect" of green finance in megacities.
Figure 6: SHAP interaction plots between GFI and TEC. The left panel shows the SHAP values of TEC across varying levels of GFI. The x-axis represents TEC, and each point is colored according to the GFI level, with lighter colors (yellow) indicating higher levels of green finance development. The y-axis represents the SHAP value of TEC, reflecting its marginal contribution to predicted CEI. The right panel provides a complementary visualization, where the x-axis shows TEC and the y-axis shows GFI. Each point represents an observation, with the color indicating the sum of their SHAP values on CEI (the contribution jointly made by the two variables). Yellow tones denote a stronger positive contribution to CEI, while purple tones indicate a negative or suppressing effect. This visualization highlights the combined marginal effects of the two features rather than their pure interaction term.
6 Conclusion
This study employs BART to investigate the complex relationships between green finance and carbon emissions across 30 Chinese provinces from 2000 to 2023. Through careful model tuning, evaluation and diagnostic testing, we confirm the strong predictive power and robustness of the BART approach in modeling environmental-economic interactions. The tuned model achieves high accuracy and generalizability, making it a reliable basis for further interpretation. Our analysis yields several key findings: (1) green finance is a relevant predictor of carbon emissions, but with a non-linear pattern. The impact of GFI on CEI exhibits a distinct inverted U-shaped pattern. In early stages of green finance development, CEI may increase due to initial resource inputs. However, once GFI surpasses a certain threshold around 0.3, it begins to significantly reduce carbon intensity. This pattern is consist with the findings of Wu et al. wu2024impact , and underscores the importance of long-term commitment and maturity in green financial systems; (2) regional heterogeneity is significant. The effects of GFI on CEI vary by region. Western regions benefit from early green finance adoption even at low levels, while Eastern provinces, where GFI is more mature. This indicates the strongest emission-reducing effects in Eastern provinces. Central China appears to be in a transition phase. These differences point to the necessity of region-specific green finance policies; (3) Green finance interacts with energy consumption, jointly influencing the outcome. Interaction analysis reveals that GFI moderates the effects of energy consumption on CEI. Specifically, high GFI can offset the emissions typically associated with high TEC, reflecting that green finance, under certain conditions, can weaken the positive correlation between energy consumption and carbon emissions, thereby playing a moderating role in emission reduction; (4) CPRI is not statistically significant. Contrary to expectations, CPRI does not exhibit a statistically significant impact on CEI in our model. However, this finding should not be interpreted as evidence that climate risk is irrelevant to carbon emissions. Rather it may stem from the limitations of using annual, province-level data, which may not adequately capture the more nuanced or delayed effects of climate risk. Climate risk often exerts its influence with a temporal lag and is likely to have a more pronounced impact at the micro level, particularly on individual firms or sectors. This highlights the need for future research that utilizes more granular data and incorporates potential lag structures to better assess the complex relationship between climate risk and emissions.
These findings suggest that while green finance is a powerful tool for reducing carbon emissions, its design and implementation must account for nonlinear effects, interaction dynamics and regional heterogeneity. In particular, in highly urbanized areas, attention should be paid to the diminishing returns of green finance, which may necessitate complementary policies, such as technological innovation or integrated urban planning. Based on these findings, the analysis can be further extended to incorporate a spatial dimension by using geographical information across provinces to examine whether spatial spillover effects or regional clustering patterns influence the effectiveness of green finance policies and carbon reduction outcomes.

### v2

@@
 Abstract
-As a core policy tool for China in addressing climate risks, green finance plays a strategically important role in shaping carbon mitigation outcomes.
 This study investigates the nonlinear and interaction effects of green finance on carbon emission intensity (CEI) using Chinese provincial panel data from 2000 to 2022.
-The Climate Physical Risk Index (CPRI) is incorporated into the analytical framework to assess its potential role in shaping carbon outcomes.
-We employ Bayesian Additive Regression Trees (BART) to capture complex nonlinear relationships and interaction pathways, and use SHapley Additive exPlanations values to enhance model interpretability.
-Results show that the Green Finance Index (GFI) has a statistically significant inverted U-shaped effect on CEI, with notable regional heterogeneity.
-Contrary to expectations, CPRI does not show a significant impact on carbon emissions.
+The Climate Physical Risk Index (CPRI) is incorporated into the analytical framework to assess the potential influence of climate risk on carbon emission performance.
+We employ Bayesian Additive Regression Trees (BART) to capture complex nonlinear relationships and interaction pathways between the variables, then use SHapley Additive exPlanations to enhance model interpretability.
+Results show that the Green Finance Index (GFI) has a significant effect on CEI, exhibiting an inverted U-shaped effect, with notable regional heterogeneity.
+Contrary to expectations, there is insufficient evidence to suggest a significant association between CPRI and CEI.
 Further analysis reveals that in high energy consumption scenarios, stronger green finance development contributes to lower CEI.
-These findings highlight the potential of green finance as an effective instrument for carbon intensity reduction, especially in energy-intensive contexts, and underscore the importance of accounting for nonlinear effects and regional disparities when designing and implementing green financial policies.
-Keywords: green finance, carbon emissions, climate risk, BART, machine learning
+These findings highlight the potential of green finance as an effective instrument for carbon reduction, especially in energy-intensive contexts, and underscore the importance of accounting for nonlinear effects and regional disparities when designing and implementing green financial policies.
+Keywords: green finance, carbon emissions, climate risk, Bayesian Additive Regression Trees
 1 Introduction
-With the escalation of global climate change and the increasing frequency of extreme weather events, carbon reduction has emerged as an urgent global priority.
+With the escalation of global climate change and the increasing frequency of extreme weather events, reducing carbon emissions has emerged as an urgent global priority.
 Green finance, as a vital strategic instrument in addressing climate change, seeks to channel financial resources into environmentally sustainable projects aimed at improving energy efficiency and reducing carbon emissions wang2023can ; wu2024impact .
-It encompasses a range of financial instruments, including green credit, green bonds, and sustainable investment funds, which effectively allocate capital to projects that mitigate the adverse impacts of climate change niyazbekova2021growth .
+Green finance encompasses a range of financial instruments, including green credit, green bonds, and sustainable investment funds, which effectively allocate capital to projects that mitigate the adverse impacts of climate change niyazbekova2021growth .
 As the world’s largest carbon emitter, China has positioned green finance as a key driver in its transition toward a low-carbon economy, particularly in pursuit of its dual carbon targets: achieving carbon peaking by 2030 and carbon neutrality by 2060 lin2023china ; du2025decarbonization .
-According to the report Fintech Facilitates Green Finance in China: Cases and Outlook Institute2023Fintech , China has become the world’s largest green credit market and the second-largest green bond market, with the scale of green finance continuing to grow.
+According to the report Institute2023Fintech , China has become the world’s largest green credit market and the second-largest green bond market, with the scale of green finance continuing to grow.
 However, the actual effectiveness of green finance is strongly shaped by local economic structures and environmental conditions.
-Recent studies suggest that climate risks significantly increase market uncertainty lee2024impact and influence investor decision-making behavior lu2025climate , thereby weakening the carbon reduction effects of green finance.
+Recent studies suggest that climate risks significantly increase market uncertainty lee2024impact and influence investor decision-making lu2025climate , thereby affecting the carbon reduction effects of green finance.
 Nevertheless, climate risk factors are rarely incorporated into analytical frameworks assessing the impact of green finance on carbon emissions.
 Moreover, most existing studies rely on traditional linear econometric models with manually specified interaction terms, and few have systematically examined the potential nonlinear effects or interaction mechanisms of green finance on carbon emissions.
-To address these research gaps, this study employs Bayesian Additive Regression Trees (BART), which are particularly suited to capturing nonlinear relationships and complex variable interactions.
+To address these research gaps, this study employs Bayesian Additive Regression Trees (BART), which is particularly suited to capture nonlinear relationships and complex variable interactions.
 Drawing on panel data covering multiple economic and environmental indicators across 30 Chinese provinces from 2000 to 2022, the analysis incorporates climate risk into the modeling framework to systematically examine how green finance affects carbon emissions and whether climate risk has a significant impact on emission outcomes.
 This research makes three key contributions to bridge the gaps in variable integration and methodological innovation: first, it incorporates climate risk into the analytical framework to assess its direct effect on carbon emissions and its potential interaction effects with green finance; second, by employing BART, it overcomes the limitations of traditional linear econometric approaches in capturing nonlinear dynamics and interaction effects.
-Combined with interpretive tools such as Partial Dependence Plots (PDPs) and SHapley Additive exPlanations (SHAP), the study reveals the underlying mechanisms through which green finance affects carbon emissions and provides interpretable empirical evidence; third, methodologically, this study expands the analytical toolkit at the intersection of green finance and carbon emissions research, demonstrating the feasibility and value of machine learning approaches in environmental policy evaluation.
-It offers technical guidance for future research and provides a robust empirical foundation for the formulation of green transition policies.
+Leveraging interpretable tools such as Partial Dependence Plots (PDPs) and SHapley Additive exPlanations (SHAP), this study reveals the underlying mechanisms through which green finance associates with carbon emissions and provides interpretable model-based evidence; third, methodologically, this study expands the analytical toolkit at the intersection of green finance and carbon emissions research, demonstrating the feasibility and value of machine learning approaches in environmental policy evaluation.
+As a consequence, this contribution offers technical guidance for future research and provides a robust empirical foundation for the formulation of green transition policies.
 2 Literature review
 2.1 Impact factors of carbon emissions
-Extensive and mature research has been conducted to identify the driving forces of carbon emissions.
-Among the various analytical frameworks, the STIRPAT model york2003stirpat remains one of the most widely applied.
-Yang et al.
+Extensive research has been conducted to identify the driving forces of carbon emissions.
+Among the various analytical frameworks, the STIRPAT model of york2003stirpat remains one of the most widely applied.
 yang2018what found a temporal shift in key emission drivers from urbanization to foreign trade.
-Tang et al.
 tang2024study further highlighted that in the metal smelting industry, population, coal consumption, and urbanization were the most sensitive factors.
-Xiao and Peng xiao2023decomposition , through Granger causality tests, confirmed that economic growth, urbanization, industrial structure, and energy structure significantly influence carbon emissions.
-To better address spatial and temporal variations, Chen et al.
-chen2021analysis adopted the GTWR-STIRPAT model and found that energy efficiency, trade, industrial structure, and environmental regulation exerted spatially heterogeneous and time-lagged effects.
-Similarly, Li & Wang li2022dynamic , using the Spatial Durbin Model and Panel Threshold Model, demonstrated that energy consumption and non-green technologies drive short-term increases in carbon emissions, whereas green innovation and industrial upgrading play a key role in long-term emission reductions.
-In addition, a number of researchers have employed decomposition approaches to examine the structural components of carbon emissions.
-Chen et al.
-chen2019driving identified several key driving factors, including potential carbon factor effects, energy intensity effects, economic scale effects, emission efficiency effects, and energy use efficiency effects.
-Liu et al.
-liu2019driving found that economic activity, emission efficiency, and potential carbon factors were the main contributors to emission growth, while improvements in energy intensity, pure technical efficiency, and scale efficiency helped mitigate emissions.
-Zhang et al.
-zhang2022carbon concluded that economic efficiency is the sole driver of industrial carbon emissions, while energy intensity remains the most critical factor in emission reduction.
-In contrast, Sun et al.
-sun2022characteristics found that energy structure and industrial structure have relatively minor effects on the growth rate of carbon emissions, and that energy intensity is the primary factor contributing to decoupling.
+Through Granger causality tests, xiao2023decomposition confirmed that economic growth, urbanization, industrial structure, and energy structure significantly influence carbon emissions.
+To better address spatial and temporal variations, chen2021analysis adopted the GTWR-STIRPAT model and found that energy efficiency, trade, industrial structure, and environmental regulation exerted spatially heterogeneous and time-lagged effects.
+Similarly, li2022dynamic , using the Spatial Durbin Model and Panel Threshold Model, demonstrated that energy consumption and non-green technologies drive short-term increases in carbon emissions, whereas green innovation and industrial upgrading play a key role in long-term emission reductions.
+In addition, several researchers have employed decomposition analysis to examine the structural components of carbon emissions.
+For examples, chen2019driving identified several key driving factors, including potential carbon factor effects, energy intensity.
+liu2019driving found that economic activity, emission efficiency, and potential carbon factors were strongly associated with emissions growth, while improvements in energy intensity, pure technical efficiency, and scale efficiency helped mitigate emissions.
+zhang2022carbon concluded that economic efficiency is the sole driver of industrial carbon emissions, while energy intensity remains the most critical factor in carbon reduction.
+In contrast, sun2022characteristics found that energy and industrial structures have relatively minor effects on the growth rate of carbon emissions, and that energy intensity is the primary factor contributing to decoupling.
 2.2 Green finance and carbon emissions
-Green finance, as an important financial tool for promoting low-carbon transition, has been widely confirmed by empirical studies to have a significant suppressive effect on carbon emissions.
-Lin et al.
+Green finance, as an important financial tool to promote the low-carbon transition, has been widely confirmed by empirical studies to have a significant suppressive effect on carbon emissions.
 lin2023impact found that green finance significantly reduces local carbon emissions and generates spillover effects in surrounding regions.
-Among various green financial instruments, Wang et al.
-wang2023can argue that the emission reduction effect of green credit is more pronounced than that of green bonds.
-Green finance not only contributes directly through investment in low-carbon projects but also facilitates carbon reduction indirectly by stimulating green technological innovation, optimizing industrial structure, and improving energy efficiency wang2022how ; tu2025impact ; ran2023driving .
-However, the emission reduction effects of green finance are not always consistent across different contexts.
+Among various green financial instruments, wang2023can argue that the carbon reduction effect of green credit is more pronounced than that of green bonds.
+Green finance not only contributes directly through investment in low-carbon projects, but also facilitates carbon reduction indirectly by stimulating green technological innovation, optimizing industrial structure, and improving energy efficiency wang2022how ; tu2025impact ; ran2023driving .
+However, the carbon reduction effects of green finance are not always consistent across different contexts.
 To explore its underlying mechanisms more deeply, recent studies have incorporated contextual variables into their analyses.
-Zhao and Li zhao2024role found that in regions with high political risk or high economic complexity, the effectiveness of green finance in reducing emissions is significantly weakened, suggesting that the institutional environment plays a moderating role.
-Ma and Fei ma2024research further emphasized that factors such as the level of governmental environmental attention and the development of the financial sector exhibit marked heterogeneity in shaping the effectiveness of green finance in influencing carbon emissions.
+zhao2024role found that in regions with high political risk or high economic complexity, the effectiveness of green finance in reducing emissions is significantly weakened, suggesting that the institutional environment plays a moderating role.
+ma2024research further emphasized that factors such as the level of governmental environmental attention and the development of the financial sector exhibit marked heterogeneity in shaping the effectiveness of green finance in influencing carbon emissions.
 Moreover, researchers have begun to explore the nonlinear nature of green finance’s impact.
-Sun sun2022characteristics and Yu et al.
-yu2024machine found that the effect of green finance on carbon emissions is not a simple linear decline.
+sun2022characteristics and yu2024machine found that the effect of green finance on carbon emissions is not a simple linear decline.
 Instead, it varies significantly under different conditions of energy consumption structure and financial technology development.
-Similarly, Cai et al.
-cai2025how uncovered complex nonlinear interactions between green finance and other factors such as environmental investment, inclusive finance, and financial regulation, all of which significantly influence carbon emission intensity.
-In summary, while the role of green finance in carbon emission reduction is increasingly evident, its underlying mechanisms exhibit notable multidimensionality, regional heterogeneity, and nonlinearity.
+Similarly, cai2025how uncovered complex nonlinear interactions between green finance and other factors such as environmental investment, inclusive finance, and financial regulation, all of which significantly influence carbon emission intensity.
+In summary, while the role of green finance in carbon emission reduction is increasingly evident, its underlying mechanisms exhibit notable complexity, regional heterogeneity, and nonlinearity.
 2.3 Application of machine learning algorithms in carbon emission
-Existing literature has made significant progress in exploring the factors influencing carbon emissions using traditional econometric models, particularly the STIRPAT framework li2017industrial ; zhang2019identifying ; thio2022estimation .
-However, these approaches generally rely on assumptions of linear relationships among variables and are often constrained by researchers’ prior selection of explanatory factors.
+The existing literature has made significant progress in exploring the factors influencing carbon emissions using traditional econometric models, particularly the STIRPAT framework li2017industrial ; zhang2019identifying ; thio2022estimation .
+However, these approaches generally rely on assumptions of linear relationships among variables and are often affected by researchers’ prior selection of explanatory factors.
 As a result, they have limitations in capturing complex nonlinear relationships and high-order interaction effects.
@@
 To address the “black-box" nature of these machine learning algorithms, some studies have further introduced approaches to improve the interpretability of results, such as SHAP (Shapley Additive Explanations) and ALE (Accumulated Local Effects).
-Gao et al.
-gao2023approach and Shan et al.
-shan2025unveiling demonstrate how these methods can reveal the marginal contributions of variables and identify nonlinear “tipping points”, thereby enhancing the interpretability of model outputs and their value for policy guidance.
+gao2023approach and shan2025unveiling , respectively, demonstrate how these methods can reveal the marginal contributions of variables and identify nonlinear “tipping points”, thereby enhancing the interpretability of model outputs and their value for policy guidance.
 In summary, the impact of green finance on carbon emissions is complex, involving multiple interacting factors cai2025how .
@@
 3 Data
-This research constructs a panel dataset that covers 30 Chinese provinces over the period 2000–2022.
-All variables are measured at the provincial level, with each variable observed separately for each province, allowing the investigation of how provincial characteristics influence carbon emission outcomes.
-A more detailed description of the 136 variables is provided in the following subsections.
-3.1 Variable description
-3.1.1 Dependent variable
-We employ carbon emission intensity (CEI) as the core indicator to measure regional carbon emission performance.
-CEI is defined as the amount of carbon emissions per unit of GDP, reflecting the carbon efficiency of regional economic activities.
-Compared to total carbon emissions, CEI controls for differences in regional economic scale, providing a more accurate measure of green development performance across regions.
+This research constructs a panel dataset covering 30 Chinese provinces over the period 2000–2022.
+All variables are measured at the provincial level, which allows us to examine how provincial characteristics influence carbon emission outcomes.
+Due to data limitations, Tibet, Hong Kong, Macau, and Taiwan are excluded.
+The following subsections provide details on the variables, their sources, and preprocessing methods.
+3.1 Dependent variable
+This study employs carbon emission intensity (CEI) as the core indicator to measure regional carbon emission performance.
+CEI is defined as the amount of carbon emissions per unit of Gross Domestic Product (GDP), reflecting the carbon efficiency of regional economic activities.
+Compared to total carbon emissions, CEI controls differences in regional economic scale, providing a more accurate measure of green development performance across regions.
 This makes it more suitable for analyzing the relationship between green finance development and regional carbon emissions, as it better reflects the environmental efficiency of economic activities under varying economic contexts.
-Following the method proposed by Wang et al. wang2022digital , this study adopts the natural logarithm of carbon emissions per unit of GDP as the response variable. The specific formula is as follows:
+Following the method proposed by wang2022digital , we adopt the natural logarithm of carbon emissions per unit of GDP as the response variable. The specific formula is as follows:
 $\mathrm{CEI}_{it}=\ln\frac{(\mathrm{CO}_{2})_{it}}{\mathrm{GDP}_{it}}.$
 where, for province $i$ in year $t$, CEI represents the logarithmic value of carbon emission intensity, $\mathrm{CO}_{2}$ denotes total carbon emissions, and GDP denotes gross domestic product.
-3.1.2 Core independent variables
-Referring to Lee and Lee LEE2022105863 , Ran et al.
-ran2023driving and Wu et al.
-wu2024impact , this study constructs a composite green finance index to measure the level of green finance development across regions.
+3.2 Core independent variable
+Referring to LEE2022105863 , ran2023driving and wu2024impact , we construct a composite green finance index as the core independent variable to measure the level of green finance development across regions.
 The index is based on seven secondary indicators: green bond, green fund, green credit, green insurance, green equity, green support, and green investment.
-These sub-indicators are first standardized to ensure cross-provincial comparability, and then aggregated using the entropy weight method to generate the final composite index.
+These sub-indicators are first standardized to ensure comparability across provinces and then aggregated using the entropy value method of ding2016 to generate the final composite index.
+The entropy value method is employed to determine the weights of the sub-indicators objectively.
+In computing the weights for the sub-indicators, the sign (positive/negative) of each indicator signal is defined following wu2024impact ; the detailed calculation procedure is provided in LEE2022105863 .
+The entropy value method reduces potential bias and ensures that the resulting green finance index more accurately reflects the underlying data characteristics.
 A higher index value indicates a higher level of green finance development in the corresponding region.
@@
 Table 1: Description of the core independent variables
-Primary indicators Secondary indicators Variable declaration
-Green Financial Index (GFI) Green bond Total green bond issuance / all bond issuance
-Green fund Market Value of Green Funds / Total Market Value of All Funds
-Green credit Credit for Environmental Protection Projects / Total Provincial Credit
-Green insurance Revenue from Environmental Liability Insurance / Total Insurance Premium Income
-Green equity Trading Volume of Carbon Emissions, Energy Use Rights, and Pollution Discharge Rights / Total Volume of Equity Market Transactions
-Green support Fiscal Expenditure on Environmental Protection / General Public Budget Expenditure
-Green investment Environmental pollution control investment / GDP
-3.1.3 Other independent variables
-Existing literature on the drivers of carbon emissions from an energy perspective has primarily focused on dimensions such as energy structure, energy efficiency, and energy intensity chen2019driving ; liu2019driving ; zhang2022carbon ; xiong2025emission .
-Meanwhile, total energy consumption (TEC) has been widely recognized as a key factor contributing to the increase in carbon emissions and has become a general consensus in the academic community li2022dynamic ; yu2024machine .
-However, when shifting the focus to carbon emission intensity, the impact of TEC becomes more complex.
-zhao2024role .
-Therefore, this paper includes TEC as one of the explanatory variables.
-On the one hand, it aims to further examine the impact of TEC on CEI.
-On the other hand, it seeks to explore whether green finance interacts with TEC in a way that facilitates the transformation of the energy consumption structure and ultimately contributes to the reduction of carbon emission intensity.
-Climate change has become a central concern globally, and managing climate risk exposure through investment strategies has emerged as a key decision-making factor for investors lu2025climate .
-Climate risk, particularly physical climate risk, can increase market uncertainty and reduce investment incentives, thereby inhibiting the development of green finance markets.
-In this context, drawing on the methodology proposed by Guo et al.
-guo2024impact , this study constructs a Climate Physical Risk Index (CPRI) based on region-specific meteorological observations and incorporates it as one of the explanatory variables in the model.
+Primary indicators Secondary indicators Weight Variable declaration
+Green Financial Index (GFI) Green bond 0.147 Total Green Bond Issuance / All Bond Issuance
+Green fund 0.144 Market Value of Green Funds / Total Market Value of All Funds
+Green credit 0.130 Credit for Environmental Protection Projects / Total Provincial Credit
+Green insurance 0.129 Revenue from Environmental Liability Insurance / Total Insurance Premium Income
+Green equity 0.148 Trading Volume of Carbon Emissions, Energy Use Rights, and Pollution Discharge Rights / Total Volume of Equity Market Transactions
+Green support 0.162 Fiscal Expenditure on Environmental Protection / General Public Budget Expenditure
+Green investment 0.140 Environmental Pollution Control Investment / GDP
+3.3 Other independent variables
+3.3.1 Total energy consumption
+Total Energy Consumption (TEC) is an important factor affecting CEI.
+Higher TEC generally increases total carbon emissions (the numerator of CEI), while greater energy use can also stimulate economic activity and expand GDP (the denominator), which may partially offset its impact on CEI.
+Including TEC as an explanatory variable allows us to examine its overall effect on CEI and explore its potential interaction with green finance in promoting a low-carbon transition.
+3.3.2 Climate physical risk index
+A higher climate physical risk may increase regional economic vulnerability to extreme weather, potentially affecting energy use, industrial output, and investment patterns, which in turn can influence CEI.
+In this context, drawing on the methodology proposed by guo2024impact , this study constructs a Climate Physical Risk Index (CPRI) based on region-specific meteorological observations and incorporates it as one of the explanatory variables in the model.
 To capture both the intensity and frequency of extreme weather events, the CPRI evaluates physical climate risk across four dimensions: extreme low temperature, extreme high temperature, extreme rainfall, and extreme drought.
 Based on daily meteorological data from local stations, the annual number of days exceeding each threshold is calculated for every province.
-To ensure comparability across provinces, each of the four sub-indices is standardized.
-The overall CPRI is then computed as the weighted average of these standardized sub-indices.
-A higher CPRI value indicates greater exposure to physical climate risk in a given region.
+In order to maintain comparability across provinces, the four sub-indices are rescaled to the same interval and then aggregated into a single value indicator.
+As a result, a higher CPRI value indicates greater exposure to physical climate risk in a given region.
+3.3.3 Socioeconomic and environmental factors
 In addition, carbon emissions are also influenced by a range of socioeconomic and environmental factors.
 Building on the existing literature from Section 2, this study incorporates several variables into the model to better capture the potential relationships among key predictors.
-These include level of economic development, industrial structure, research and development (R&D) intensity, urbanization level, population density, foreign direct investment, degree of government intervention, environmental regulation intensity.
+These include the level of economic development, industrial structure, research and development (R&D) intensity, urbanization level, population density, foreign direct investment, degree of government intervention, environmental regulation intensity.
 In light of empirical studies that have confirmed the significant regional heterogeneity in the impact of green finance on carbon emissions, this study further divides the provinces into eastern, central, and western regions based on the traditional regional classification method, and incorporates this classification as a categorical variable in the model.
 3.4 Data sources
-This study constructs a panel dataset covering 30 provinces in China from 2000 to 2022.
-Due to severe data limitations, the Tibet, Hong Kong, Marco and Taiwan is excluded from the analysis.
-A more detailed description of the data is shown in Table 2.
+All data are obtained from official sources, including the China Statistical Yearbook, China Energy Statistical Yearbook, Regional Statistical Yearbooks, CSMAR database, Wind database, the Environmental Status Bulletin, and the China Deep Data database.
+Table 2 provides a detailed description of all variables and sources.
 Table 2: Data description
@@
 3.5 Data preprocessing
-To reduce right-skewness in the distribution of variables, we applied logarithmic transformations to the dependent variable (CEI) and to selected independent variables, namely economic development (GDP), population density (population), industrial structure (industry), and environmental regulation (environment), all of which exhibited skewness statistics greater than 2.
+To reduce right-skewness in the distribution of variables, we applied logarithmic transformations to the dependent variable (CEI) and to the selected independent variables, namely economic development (GDP), population density (population), industrial structure (industry), and environmental regulation (environment), all of which exhibited a skewness value greater than 2.
 We then employed Generalized Variance Inflation Factors (GVIF) to conduct a preliminary analysis of potential multicollinearity among explanatory variables.
 A GVIF value close to 1 indicates little to no multicollinearity, while values equal to or greater than 5 or 10 suggest strong multicollinearity obrien2007caution .
-Similarly, the adjusted generalized standard error inflation factor (aGSIF) values exceeding 2.2 or 3.2 are also considered indicative of significant multicollinearity (Fannie et al., 2025).
+Similarly, we consider the adjusted generalized standard error inflation factor (aGSIF) for which values exceeding 2.2 or 3.2 are also considered indicative of significant multicollinearity shabangu2025monthly .
 As shown in Table 3, the GVIF for the variable representing the level of economic development is 5.55 (>>5), and the corresponding aGSIF is 2.36 (>>2.2), indicating notable multicollinearity.
@@
 4.1 Model specification
-Originally proposed by Chipman et al.
-chipman2010bart , Bayesian Additive Regression Trees (BART) is a flexible, nonparametric Bayesian ensemble method that represents the response variable $y$ as the sum of $m$ regression trees:
+Originally proposed by chipman2010bart , Bayesian Additive Regression Trees (BART) is a flexible, nonparametric Bayesian ensemble method that represents the response variable $y$ as the sum of $m$ regression trees:
 $y=\sum_{j=1}^{m}g(x;T_{i},M_{i})+\varepsilon,\quad\varepsilon\sim N(0,\sigma^{2}),$
@@
 This is achieved by computing the partial residual:
-$R_{i}=y-\sum_{k\neq i}g(x;T_{k},M_{k})$
+$R_{i}=y-\sum_{k\neq i}g(x;T_{k},M_{k}),$
 which represents the part of the response not yet explained by the other trees.
@@
 By fitting each tree to the current residuals, the algorithm gradually refines the overall model, effectively adapting to nonlinear patterns and latent interaction effects in the data.
-The update of the tree structure $T_{i}$ is conducted using a Metropolis–Hastings (MH) step.
+The update of the tree structure $T_{i}$ is conducted using a Metropolis–Hastings step.
 New tree proposals are generated through operations such as growing or pruning nodes, modifying split rules, or swapping decision nodes.
@@
 The noise variance $\sigma^{2}$ is updated at the end of each iteration by drawing from a full conditional inverse-gamma distribution.
-By running the MCMC algorithm for a sufficiently large number of iterations and discarding the initial samples as burn-in, a sequence of fitted functions is obtained:
+The MCMC algorithm was run for a large number of iterations, discarding the initial samples as burn-in.
+Satisfactory convergence was checked using trace plots and standard convergence diagnostics.
+From the resulting posterior samples, a sequence of fitted functions is obtained:
 $f^{(k)}(x)=\sum_{i=1}^{m}g(x;T_{i}^{(k)},M_{i}^{(k)}).$
@@
 The model framework and sampling algorithm together enable BART to flexibly capture complex nonlinear relationships and high-order interactions, even when such interactions are not explicitly specified in the model.
-The model is evaluated using Root Mean Square Error (RMSE) and pseudo-$R^{2}$. A lower RMSE indicates a better model fit. The RMSE is calculated as follows:
+The model fit is evaluated using Root Mean Square Error (RMSE) and pseudo-$R^{2}$. A lower RMSE indicates a better fit to the data. The RMSE is calculated as follows:
 $RMSE=\sqrt{\frac{1}{n}\sum_{i=1}^{n}(y_{i}-\hat{y}_{i})^{2}},$
-The pseudo-$R^{2}$ is calculated using the formula:
+whereas the pseudo-$R^{2}$ is calculated using the formula:
 $\text{pseudo-}R^{2}=1-\frac{SSE}{SST}.$
-where SSE is the sum of squared errors and SST is the total sum of squares.
-It indicates the proportion of variation in the dependent variable that is explained by the model.
+Here, SSE is the sum of squared errors and SST is the total sum of squares, indicating the proportion of variation in the dependent variable that is explained by the model.
 A higher pseudo-$R^{2}$ suggests better explanatory power.
-4.2 Interpretability methods
-Despite the strong predictive performance of the BART model in both regression and classification tasks, it is often considered a “black-box" predictor due to its lack of transparency and interpretability chipman2010bart .
-Kapelner and Bleich kapelner2016bartmachine incorporated visualization tools such as partial dependence plots (PDPs) into the bartMachine framework.
+4.2 Summary tools
+Despite the strong predictive performance of the BART model in both regression and classification tasks, As chipman2010bart mentioned, it is often considered a “black-box" predictor due to its complex model structure which cannot be easily summarized.
+kapelner2016bartmachine incorporated visualization tools such as partial dependence plots (PDPs) into the bartMachine framework.
 However, while PDPs provide insight into average marginal effects, they may fall short in capturing more complex or localized feature interactions.
-To complement PDPs and provide a more nuanced interpretation of model behavior, this study also employs SHapley Additive exPlanations (SHAP).
-These tools examine feature contributions from different theoretical perspectives, thereby enhancing the depth and reliability of the model interpretation, especially for complex interactions and heterogeneous effects.
+To complement PDPs and provide a more nuanced interpretation of model behavior, this study also employs SHapley Additive exPlanations lundberg2017unified .
+These tools examine feature contributions from different perspectives, thereby enhancing the depth and reliability of the model interpretation, especially for complex interactions and heterogeneous effects.
 4.2.1 Partial dependence plot
-This study employs the Partial Dependence Plot (PDP) from the bartMachine framework to examine how a specific predictor affects the response variable on average, while controlling for other predictors.
-The PDP is based on the formulation introduced by Friedman friedman2001greedy :
+We adopt PDPs from the bartMachine framework to examine how a specific predictor affects the response variable on average, while controlling for other predictors.
+The PDP is based on the formulation introduced by friedman2001greedy :
 $f_{j}(x_{j})=E_{x_{-j}}[f(x_{j},x_{-j})]:=\int f(x_{j},x_{-j})dP(x_{-j}),$
+where $f$ represents the true but unknown model function, and $dP(x_{-j})$ denotes the marginal distribution of all covariates except the jthj-th one.
+The function $f_{j}$ measures the marginal or average effect of a single predictor $x_{j}$ on the response variable.
+Its principle is to “average out” the influence of the other variables by integrating over their marginal distribution.
+In other words, $f_{j}(x_{j})$ describes how the expected model prediction changes as $x_{j}$ varies, while the effects of the other variables are held at their average levels, thereby isolating the independent contribution of a single predictor to the model’s output.
 Since the true model function $f$ and the marginal distribution $dP(x_{-j})$ are unknown, this expectation is approximated by averaging over the training data:
-$\hat{f}_{j}(x_{j})=\frac{1}{n}\sum_{i=1}^{n}\hat{f}(x_{j},x_{-j,i})$
-where $x_{-j,i}$ represents all variables except $x_{j}$ from the ithi-th training sample, and $\hat{f}$ denotes predictions generated by the bartMachine model.
+$\hat{f}_{j}(x_{j})=\frac{1}{n}\sum_{i=1}^{n}\hat{f}(x_{j},x_{-j,i}).$
 where $x_{-j,i}$ represents all variables except $x_{j}$ from the ithi-th training sample, and $\hat{f}$ denotes predictions generated by the bartMachine model.
@@
 4.2.2 SHapley Additive exPlanations
-SHAP (SHapley Additive exPlanations) is grounded in the game-theoretic concept of Shapley values, which were originally proposed for model explanation by Štrumbelj & Kononenko strumbelj2011general ; strumbelj2014explaining .
-However, the approach only gained widespread attention after Lundberg & Lee lundberg2017unified introduced SHAP as a unified framework for feature attribution.
-Their work not only rebranded the method but also extended its theoretical foundations, connecting Shapley values with other model-agnostic interpretability techniques such as LIME.
+SHAP (SHapley Additive exPlanations) is a method grounded in the game-theoretic concept of Shapley values, which were originally proposed for model explanation by strumbelj2011general ; strumbelj2014explaining .
+However, the approach only gained widespread attention after lundberg2017unified introduced SHAP as a unified framework for feature attribution.
+Their work not only rebranded the method but also extended its theoretical foundations, connecting Shapley values with other model-agnostic interpretability techniques such as LIME ribeiro2016why .
 SHAP treats features as “players" in a cooperative game, where the model prediction is the “payout" to be fairly distributed among them.
@@
 Formally, the Shapley value for feature $j$, given a prediction function $f$ and input $x$, is defined as:
-$\phi_{i}=\sum_{S\subseteq F\setminus\{i\}}\frac{|S|!\cdot(|F|-|S|-1)!}{|F|!}[f(S\cup i)-f(S)]$
-where $F$ is the set of all features; $S$ is the subset of features excluding $i$; $f(S)$ is the expected prediction when only features in $S$ are known.
+$\phi_{i}=\sum_{S\subseteq V\setminus\{i\}}\frac{|S|!\cdot(|V|-|S|-1)!}{|V|!}[f(S\cup i)-f(S)].$
+where $V$ is the set of all features; $S$ is the subset of features excluding $i$; $f(S)$ is the expected prediction when only features in $S$ are known.
 In practice, for models with a large number of input variables, exhaustively enumerating all feature subsets S to compute Shapley values is computationally infeasible.
-To address this challenge, we employed the FastSHAP R package developed by Greenwell and Greenwell greenwell2020fastshap , which provides an efficient approximation of Shapley values for any supervised learning model.
-Rather than calculating exact Shapley values by iterating over all possible feature coalitions, FastSHAP employs a Monte Carlo-based approximation method proposed by Strumbel & Kononenko strumbelj2014explaining .
+To address this challenge, we employed the FastSHAP R package developed by greenwell2020fastshap , which provides an efficient approximation of Shapley values for any supervised learning model.
+Rather than calculating exact Shapley values by iterating over all possible feature coalitions, FastSHAP employs the Monte Carlo-based approximation method proposed by strumbelj2014explaining .
 The approximated outputs, commonly referred to as SHAP values, preserve the theoretical properties of Shapley values while enabling scalable interpretation of complex black-box models.
 In our analytical framework, we first use Partial Dependence Plots (PDPs) from the bart package to explore the global structure and interaction patterns of the model.
-Then, we apply SHAP values to conduct detailed mechanism analysis.
-This combined approach helps provide a more comprehensive understanding of how input features affect the predictions and enhances the transparency and credibility of BART.
+Then, we complement PDPs with SHAP values to conduct a more in depth analysis of the interactions between variables analysis.
+This combined approach helps to provide a more comprehensive understanding of how input features affect the predictions, thus enhancing the transparency and reliability of BART.
 5 Results
 5.1 Model performance
-5.1.1 Model evaluation
 To ensure a realistic evaluation of the model’s predictive performance and to avoid data leakage, we used data from 2000 to 2020 as the training set and data from 2021 to 2022 as the test set.
 Additionally, we applied 10-fold cross-validation to evaluate the model’s generalization ability.
-The BART model initially trained with default hyperparameters achieved a pseudo-R² of 0.989 and an RMSE of 0.1 on the training set.
+The BART model initially trained with default hyperparameters achieved a pseudo-R² of 0.989 and an RMSE of 0.1 in the training set.
 The 10-fold cross-validation yielded a pseudo-R² of 0.9586 and an RMSE of 0.1886.
-On the test set, the model achieved a pseudo-R² of 0.8493 and an RMSE of 0.2826.
-These results indicate that even with default settings, BART demonstrates strong fitting ability and successfully captures most of the underlying data patterns.
-However, the model’s predictive performance on unseen data was relatively weaker, and the residuals failed the normality test (p-value=0.00091<0.05), indicating potential model misspecification or non-Gaussian error distribution.
+In the test set, the model achieved a pseudo-R² of 0.8493 and an RMSE of 0.2826.
+These results indicate that even with default settings, BART achieves a good fit strong fitting ability and successfully captures most of the underlying data patterns.
+However, the model’s predictive performance on unseen data was relatively weaker, and the residuals failed the normality test (p-value=0.00091 << 0.05), indicating potential model misspecification or non-Gaussian error distribution.
 To address this, we used the built-in grid search function in the bartMachine package to tune the model’s hyperparameters.
@@
 Moreover, the residuals from the tuned model passed the normality test (p-value = 0.269 >> 0.05), indicating a better fit to the data.
-In summary, the tuned BART model demonstrates both high predictive power and robust generalization.
+In summary, the tuned BART model seems to achieve both high predictive power and robust generalization.
 Therefore, all subsequent analyses on the relationships between green finance, climate risk, and carbon emission intensity are based on the tuned model.
-5.1.2 Model Diagnostics and Statistical Significance Testing
+5.1.1 Preliminary model diagnostics
 Figure 1 presents the residuals Q-Q plot to assess normality and the residuals vs predicted values plot to assess homoskedasticity.
-The Shapiro-Wilk test for residuals yielded a p-value of 0.269 >> 0.05, indicating that the residual distribution generally conforms to normality.
+The Shapiro-Wilk test for residuals yielded a p-value of 0.269 >> 0.05, indicating that the assumption of normality should not be rejected.
 Although a few points deviate in the tails of the Q-Q plot, the majority of points align closely along the diagonal line.
-Meanwhile, the residuals vs predicted values plot shows no obvious pattern or trend, suggesting that there is no significant heteroskedasticity issue.
+Meanwhile, the residuals vs predicted values plot shows no obvious pattern or trend, suggesting that there are no particular concerns regarding heteroskedasticity.
+Figure 1: Test of normality of errors using $Q$-plot and the Shapiro-Wilk test (left), residual plot to assess heteroskedasticity (right).
+5.2 Feature importance and interaction terms
 In addition, bartMachine provides a permutation-based approach to assess the statistical significance of individual predictors or groups of variables, analogous to the partial F-test in ordinary least squares regression but without assuming linearity.
 We applied this method to test the overall model validity, as well as the individual significance of the GFI and CPRI.
-For the overall model test, we permuted the response variable $y$, thereby breaking any association between the predictors and the outcome.
+For the overall model test, we shuffled the order of the response variable $y$, thereby breaking any association between the predictors and the outcome.
 This is equivalent to an omnibus F-test in linear regression.
@@
 In each case, the values of the target variable were randomly permuted within the sample, destroying its relationship with the response while keeping all other variables fixed.
-This allows us to test whether the variable contributes predictive power after controlling for others.
-The results show that GFI has a p-value of 0.0396 < 0.05, indicating a statistically significant effect on carbon emission intensity.
+This allows us to test whether the variable contributes predictive power after controlling for the others.
+The results show that GFI has a p-value of 0.0396 << 0.05, indicating a statistically significant effect on carbon emission intensity.
 In contrast, CPRI yields a p-value of 0.238, which exceeds the 0.05 threshold.
 Therefore, we fail to reject the null hypothesis and conclude that there is insufficient statistical evidence to support a significant effect of CPRI on carbon emission intensity.
-Also, we conducted individual significance tests for GFI and CPRI.
-In each case, the values of the target variable were randomly permuted within the sample, destroying its relationship with the response while keeping all other variables fixed.
-This allows us to test whether the variable contributes predictive power after controlling for others.
-The results show that GFI has a p-value of 0.02, indicating a statistically significant effect on carbon emission intensity.
-In contrast, CPRI yields a p-value of 0.281, which exceeds the 0.05 threshold.
-Therefore, we fail to reject the null hypothesis and conclude that there is insufficient statistical evidence to support a significant effect of CPRI on carbon emission intensity.
-Figure 1: Test of normality of errors using $Q$-plot and the Shapiro-Wilk test (left), residual plot to assess heteroskedasticity (right).
 5.3 Feature importance and interaction terms
-This section utilizes the built-in functionalities of bartMachine to examine the importance of individual predictors and their interactions in BART, Additionally, SHAP methods are employed to visualize nonlinear relationships and interactions among variables.
-Figures 2, 3, and 4 present the variable importance and interaction results extracted from bartMachine, while Figures 5, and 6 show visualizations based on SHAP.
-As shown in Figure 2, TEC appears most frequently in the model’s decision paths and has relatively narrow 95% credible intervals (CI), indicating that it contributes the most to explaining variations in CEI, with a high degree of certainty.
-GFI ranks fifth, and its corresponding 95% CI is also relatively narrow, suggesting that green finance exerts a significant and stable influence within the model.
-In contrast, CPRI ranks the lowest, with a near-zero contribution, implying limited predictive relevance.
-The statistical significance of each variable are further validated in Section 5.1.2.
-Regarding interaction effects, Figure 3 displays the top 10 variable pairs with the strongest interactions.
-Among them, GFI is involved in several interactions, including GFI × TEC, GFI × population, and GFI × govern.
-This suggests that the impact of green finance on CEI largely depends on its interaction with other key variables.
-Notably, the GFI × TEC interaction ranks fifth overall and is the most prominent among GFI-related pairs.
-Therefore, the subsequent analysis will focus on the interaction between GFI and TEC to explore the mechanism through which green finance influences the marginal effect of energy consumption on carbon emission intensity.
+This section utilizes the built-in functionalities of bartMachine to examine the importance of individual predictors and their interactions within BART.
+In the BART framework, variable importance is evaluated using the inclusion proportion metric chipman2010bart ; bleich2014variable .
+The inclusion proportion for a given predictor represents the frequency with which that variable is selected as a splitting rule across all posterior samples (among all tree structures generated during the MCMC iterations).
+A higher inclusion proportion indicates that the variable appears more frequently in the model’s partitioning process and thus contributes more substantially to explaining variations in the response variable.
+Interaction effects are evaluated following the approach of damien2013bayesian , where an interaction between two predictors is considered present when both variables appear within splitting rules along a path in the same tree.
+To quantify these interactions, the method counts how often each pair of variables co-occurs in such splits across all trees and posterior samples.
+Summing these counts provides an overall measure of the interaction strength for each pair, which can then be used to assess the relative importance of their combined effect on the response.
+This approach allows the identification of predictor pairs whose joint influence is substantial, including cases where a variable may interact with itself.
 Figure 2: The average inclusion proportions of explanatory variables in the tuned BART model, reflecting how frequently each variable appears in the model’s decision paths.
@@
 The segments atop the bars indicate the 95% confidence intervals.
+As shown in Figure 2, TEC appears most frequently in the model’s decision paths and has relatively narrow 95% credible intervals, indicating that it contributes the most to explaining variations in CEI, with a high degree of certainty.
+GFI ranks fifth, and its 95% credible interval is relatively wide compared with those of other predictors, suggesting that although green finance exerts a significant influence, the model exhibits some uncertainty about the precise magnitude of its effect.
+In contrast, CPRI ranks the lowest, with a near-zero inclusion proportion and one of the narrowest 95% credible interval, suggesting that it plays a limited role in explaining variations in CEI and that the model is fairly certain about this result.
+The statistical significance of each variable are further validated in Section 5.1.1.
+In terms of interaction effects, bartMachine constructs 25 independent BART models, each trained from a different random initialization (randomly chosen initial tree structure and parameter state).
+The interaction strengths are then averaged across these models to obtain stable estimates that account for different posterior modes in the tree and distribution.
+Figure 3 displays the top 10 variable pairs with the strongest interactions, as measured by their average co-occurrence across decision paths in 25 constructed models.
+Among these, GFI appears frequently as a component of key interacting pairs, including GFI × TEC, GFI × population, and GFI × govern.
+The presence of GFI in multiple high-ranking interactions indicates that its influence on CEI may be highly conditional on other structural and socioeconomic factors.
+In particular, the interaction between GFI and TEC ranks fifth overall and stands out as the most influential GFI-related pair.
+This finding suggests that the marginal impact of green finance on carbon emissions may be amplified or moderated depending on the level of energy consumption.
+Therefore, the subsequent analysis will focus on the interaction between GFI and TEC to explore the mechanism through which green finance influences the marginal effect of energy consumption on carbon emission intensity.
 Figure 3: In the tuned BART model, this figure displays the top 10 variable pairs ranked by average interaction counts, based on 25 model constructions.
@@
 5.4 Nonlinear and interaction effects analysis
-5.3.1 Nonlinear and regional effects of GFI
 Figures 4 and 5 jointly explore the nonlinear and heterogeneous effects of GFI on CEI, using PDP from BART and SHAP visualizations.
-Figure 4 displays the PDP of GFI, illustrating the average marginal effect of GFI on CEI while holding all other variables constant.
-The results reveal a clear inverted U-shaped pattern: at low levels of GFI, the PDP curve rises, indicating that the initial development of green finance is associated with an increase in CEI.
-However, once GFI exceeds a certain threshold (around 0.3), the curve begins to decline, suggesting that further development of green finance helps reduce CEI.
-This implies that green finance has greater emission reduction potential in more advanced stages.
-The relatively narrow 95% confidence interval further confirms the stability and robustness of this nonlinear relationship.
-The SHAP-based analysis in Figure 5 further reveals regional heterogeneity in the observed nonlinear relationship.
-In Eastern and Central China, the inverted U-shape is more pronounced, with most observations concentrated within the GFI range of 0.2-0.4.
-After GFI exceeds 0.3, SHAP values in both regions show a general downward trend.
-However, in Central China, most SHAP values remain positive, indicating that although GFI has reached a relatively high level, its emission reduction effect has not yet fully materialized, possibly due to the green finance sector still being in its expansion or investment phase.
-In contrast, in Eastern China, SHAP values decline more significantly after GFI exceeds 0.3, with some observations even turning negative.
-This may reflect a more mature green finance system in the region, which is beginning to effectively curb carbon intensity.
-In Western China, GFI levels are more widely distributed.
-When GFI is relatively low, SHAP values are mostly negative, suggesting that even in its early stages, green finance has already contributed to reducing CEI to some extent.
-As GFI increases, SHAP values gradually rise and approach positive territory, indicating that the emission reduction effect weakens during the mid-development stage.
-However, when GFI reaches approximately 0.3, SHAP values begin to decline again, showing signs of diminishing marginal returns or a “saturation effect.”
-These findings confirm that the effectiveness of green finance in reducing carbon emissions is nonlinear and regionally differentiated, highlighting the importance of region-specific strategies in promoting green finance development.
+The PDP curve in Figure 4 exhibits a typical inverted U-shaped pattern: at lower GFI levels, CEI increases with GFI, but once GFI exceeds a threshold of approximately 0.3, the curve begins to decline, indicating an emission-reduction effect at higher levels of GFI.
+However, the interpretability of this nonlinear pattern depends on the underlying data-generating process.
 Figure 4: PDP of GFI on CEI.
@@
 The narrowing of the credible interval suggests increasing model certainty.
-The x-axis shows GFI values at specified quantiles (5%, 10%, …, 95%), while the y-axis represents the average predicted CEI at each GFI level, conditional on other variables being fixed.
-Figure 5: SHAP PDP of GFI on CEI, color-coded by region.
+The x-axis shows GFI values, while the y-axis represents the average predicted CEI at each GFI level, conditional on other variables being fixed.
+Figure 5: SHAP PDP of GFI on CEI, color-coded by region (red for Eastern, purple for Central, yellow for Western).
 Each point represents an observation, with the x-axis showing the value of GFI and the y-axis showing the corresponding SHAP value, which reflects the marginal contribution of GFI to the model prediction of CEI.
@@
 A positive SHAP value indicates that GFI increases predicted CEI (i.e., contributes positively), while a negative value suggests a suppressing effect (i.e., contributes negatively).
-The color of the points denotes regional classification: Eastern (red), Central (purple), and Western (yellow).
+The PDP is constructed by averaging over the effects of other variables while allowing GFI to vary, which implicitly assumes that GFI is independent or only weakly correlated with other covariates.
+In real economic systems, this assumption rarely holds, particularly in the low-GFI range where green finance is often closely related to factors such as technological progress, industrial structure, and policy implementation.
+Therefore, the apparent “positive correlation” observed at low GFI levels is more likely to reflect unobserved confounding effects rather than the true marginal impact of GFI.
+For example, in regions with lower technological development or less optimized economic structures, both insufficient green finance and higher carbon emissions tend to occur simultaneously, resulting in a spurious positive relationship in the model.
+Once GFI exceeds the threshold of approximately 0.3, the influence of these confounding factors gradually weakens, and the independent role of green finance begins to emerge.
+The downward trend in this range more likely reflects the genuine emission-reduction effect of green finance.
+On one hand, a higher level of GFI channels financial and policy resources toward low-carbon, energy-efficient, and renewable energy projects, directly lowering carbon emission intensity.
+On the other hand, the expansion of green finance may also stimulate technological innovation and green industrial upgrading, establishing a long-term mitigation mechanism.
+In addition, a mature green finance system can enhance the environmental governance capacity of enterprises and local governments, leading to more efficient energy use and emission control.
+Therefore, the most meaningful part of Figure 4 lies in the high-GFI range, where green finance demonstrates a tangible and economically interpretable marginal effect on reducing CEI.
+This indicates that the environmental benefits of green finance become effective only when it reaches a sufficiently advanced level of development.
+Figure 5 presents the SHAP-based analysis, which further reveals the regional heterogeneity underlying the nonlinear relationship observed in the PDP results.
+As shown in the figure, most samples in the range of GFI below approximately 0.3 are concentrated in the central and western regions, whereas those from the eastern region are primarily distributed at higher GFI levels above 0.3.
+It is important to note that although the SHAP values for the western region are generally negative at low GFI levels, this does not imply that low GFI exerts a genuine emission-reduction effect.
+As discussed in the above, low GFI levels are often accompanied by limited technological innovation capacity, slower industrial upgrading, and weaker policy implementation.
+Therefore, the relationship observed in this range likely reflects unobserved structural factors or confounding effects rather than the true marginal contribution of GFI.
+Beyond the threshold of approximately 0.3, the carbon reduction effect of GFI becomes evident and displays notable regional heterogeneity.
+In the central and western regions, SHAP values exhibit an overall downward trend but remain mostly positive, suggesting that although green finance has developed to a higher level, its carbon reduction potential has not yet been fully realized.
+These regions may still be in a stage dominated by investment expansion and institutional development, where economic foundations, technological capabilities, and policy frameworks require further improvement.
+In contrast, SHAP values in the eastern region show a more pronounced decline at higher GFI levels, with some observations even turning negative.
+This indicates that in regions with stronger economic foundations, greater technological innovation, and more mature policy systems, advanced green finance development has begun to exert a substantial and measurable mitigating effect on carbon emission intensity, reflecting the effectiveness of a well-established and efficient green finance system.
+Taken together with the PDP results, these findings demonstrate that the carbon reduction effects of green finance are not only nonlinear but also regionally heterogeneous.
+More importantly, the genuine marginal impact of green finance tends to emerge only once it surpasses a certain development threshold, and both the strength and timing of this effect vary significantly across regions.
+This evidence underscores the need for region-specific strategies that align with local economic structures and development stages to promote green finance as a sustained mechanism for reducing carbon intensity.
 5.4.1 Interaction effects between GFI and TEC
-Figure 6 presents SHAP interaction effect plots between TEC and GFI, aiming to gain insights into complex feature interactions and uncover latent patterns.
-The left panel shows that when TEC is relatively low (below 10000) and GFI is also at a low level, the SHAP values are predominantly positive.
-This indicates that under conditions of low energy consumption and limited green finance development, carbon emission intensity tends to increase.
-As TEC increases beyond this level, accompanied by rising GFI levels, the SHAP values exhibit a clear downward trend and eventually turn negative.
-This suggests that when energy consumption is high, the interaction with higher levels of green finance leads to a negative marginal effect on CEI.
-In other words, the expansion of green finance may help decouple energy consumption from emissions, alleviating the environmental pressure typically associated with high energy use.
-In the right panel, as both TEC and GFI increase, the SHAP values shift from yellow to purple.
-This color gradient visually reinforces the finding that the marginal harm of energy consumption on carbon intensity becomes weaker, or even beneficial under high GFI conditions.
-Taken together, the interaction between GFI and TEC reveals a clear nonlinear and synergistic relationship in shaping carbon emission intensity.
-Green finance not only directly influences CEI but also plays a critical moderating role by mitigating the environmental cost of energy use.
+Figure 6 presents the SHAP interaction effects between total energy consumption (TEC) and green finance (GFI), illustrating how their joint variations influence carbon emission intensity (CEI).
+It can be observed that when both TEC and GFI are relatively low, the SHAP values are predominantly positive, indicating that limited energy use and underdeveloped green finance are associated with higher CEI levels.
+As TEC and GFI increase, the SHAP values exhibit a downward trend, eventually even turning negative.
+It should be noted, however, that this decline in SHAP values does not imply that higher energy consumption directly reduces carbon intensity.
+This phenomenon can be explained by the definition of CEI and the potential role of green finance.
+In this study, CEI is defined as carbon emissions per unit of GDP, and higher TEC is generally accompanied by increased economic output, which expands the GDP denominator and leads to a lower measured CEI.
+High levels of green finance may promote improvements in energy efficiency and transitions to cleaner energy sources, thereby weakening the marginal impact of energy consumption on CEI.
+Moreover, in more advanced stages of green finance development, economic scale expansion may offset part of the emission-promoting effect of increased energy consumption, further reducing carbon intensity per unit of output.
+In other words, the observed decline in SHAP values reflects the combined moderating effects of green finance and economic growth on carbon intensity, rather than a direct emission-reducing effect of energy consumption.
+Figure 6: SHAP interaction plots between GFI and TEC.
+The plot shows the SHAP values of TEC across varying levels of GFI.
+The x-axis represents TEC, and each point is colored according to the GFI level, with lighter colors (yellow) indicating higher levels of green finance development.
+The y-axis represents the SHAP value of TEC, reflecting its marginal contribution to predicted CEI.
+This visualization highlights the combined marginal effects of the two features rather than their pure interaction term.
+Overall, the interaction between GFI and TEC reveals a clear nonlinear and synergistic relationship in shaping carbon emission intensity.
+Specifically, green finance not only directly influences CEI but also plays a critical moderating role by weakening the marginal effect of energy consumption on carbon intensity.
 These findings emphasize the importance of integrating financial instruments into carbon mitigation strategies.
 In particular, energy-intensive regions stand to benefit more from targeted green finance policies that can curb emissions even amid rising energy demand.
-Moreover, accounting for such interaction effects contributes to a more nuanced understanding of how various policy instruments, including financial incentives, energy regulation and regional planning, can be effectively coordinated to achieve long-term environmental objectives.
-In summary, the interaction between population and GFI in shaping carbon emission intensity is highly complex and stage-dependent.
-These findings highlight the need for regionally differentiated green finance policies and call for particular attention to the potential “saturation effect" of green finance in megacities.
-Figure 6: SHAP interaction plots between GFI and TEC.
-The left panel shows the SHAP values of TEC across varying levels of GFI.
-The x-axis represents TEC, and each point is colored according to the GFI level, with lighter colors (yellow) indicating higher levels of green finance development.
-The y-axis represents the SHAP value of TEC, reflecting its marginal contribution to predicted CEI.
-The right panel provides a complementary visualization, where the x-axis shows TEC and the y-axis shows GFI.
-Each point represents an observation, with the color indicating the sum of their SHAP values on CEI (the contribution jointly made by the two variables).
-Yellow tones denote a stronger positive contribution to CEI, while purple tones indicate a negative or suppressing effect.
-This visualization highlights the combined marginal effects of the two features rather than their pure interaction term.
+Moreover, accounting for such interaction effects, including TEC × GFI, contributes to a more nuanced understanding of how various policy instruments, such as financial incentives, energy regulation, and regional planning, can be effectively coordinated to achieve long-term environmental objectives.
 6 Conclusion
-This study employs BART to investigate the complex relationships between green finance and carbon emissions across 30 Chinese provinces from 2000 to 2023.
+This study employs BART to investigate the complex relationships between green finance and carbon emissions across 30 Chinese provinces from 2000 to 2022.
 Through careful model tuning, evaluation and diagnostic testing, we confirm the strong predictive power and robustness of the BART approach in modeling environmental-economic interactions.
@@
 The impact of GFI on CEI exhibits a distinct inverted U-shaped pattern.
-In early stages of green finance development, CEI may increase due to initial resource inputs.
-However, once GFI surpasses a certain threshold around 0.3, it begins to significantly reduce carbon intensity.
-This pattern is consist with the findings of Wu et al.
-wu2024impact , and underscores the importance of long-term commitment and maturity in green financial systems; (2) regional heterogeneity is significant.
-The effects of GFI on CEI vary by region.
-Western regions benefit from early green finance adoption even at low levels, while Eastern provinces, where GFI is more mature.
-This indicates the strongest emission-reducing effects in Eastern provinces.
-Central China appears to be in a transition phase.
-These differences point to the necessity of region-specific green finance policies; (3) Green finance interacts with energy consumption, jointly influencing the outcome.
-Interaction analysis reveals that GFI moderates the effects of energy consumption on CEI.
-Specifically, high GFI can offset the emissions typically associated with high TEC, reflecting that green finance, under certain conditions, can weaken the positive correlation between energy consumption and carbon emissions, thereby playing a moderating role in emission reduction; (4) CPRI is not statistically significant.
+This result generally aligns with the previous findings of wu2024impact .
+At low levels of GFI, the observed positive relationship with CEI likely reflects confounding structural factors rather than a direct effect of green finance.
+Once GFI surpasses a threshold of approximately 0.3, its genuine emission-reduction effect emerges, highlighting the importance of advanced development and long-term commitment in green financial systems.
+(2) Regional heterogeneity is apparent.
+The effects of GFI on CEI vary across regions.
+Eastern provinces, where green finance has a stronger presence, show the most pronounced carbon reduction effects at higher GFI levels.
+Central and Western regions, in contrast, remain in a transitional stage, with the mitigation potential of green finance not yet fully realized.
+These differences emphasize the need for region-specific strategies that account for local economic structures, technological capabilities, and policy maturity.
+(3) Green finance interacts with energy consumption, jointly influencing the outcome.
+GFI moderates the effects of TEC on CEI.
+High GFI can offset the positive emissions effects typically associated with high TEC, reflecting that green finance can weaken the correlation between energy consumption and carbon emissions under certain conditions.
+This underscores the importance of integrating financial instruments into carbon mitigation strategies, particularly in energy-intensive regions; (4) CPRI is not statistically significant.
 Contrary to expectations, CPRI does not exhibit a statistically significant impact on CEI in our model.
 However, this finding should not be interpreted as evidence that climate risk is irrelevant to carbon emissions.
-Rather it may stem from the limitations of using annual, province-level data, which may not adequately capture the more nuanced or delayed effects of climate risk.
+Rather, it may stem from the limitations of using annual, province-level data, which may not adequately capture the more nuanced or delayed effects of climate risk.
 Climate risk often exerts its influence with a temporal lag and is likely to have a more pronounced impact at the micro level, particularly on individual firms or sectors.
@@
 These findings suggest that while green finance is a powerful tool for reducing carbon emissions, its design and implementation must account for nonlinear effects, interaction dynamics and regional heterogeneity.
-In particular, in highly urbanized areas, attention should be paid to the diminishing returns of green finance, which may necessitate complementary policies, such as technological innovation or integrated urban planning.
 Based on these findings, the analysis can be further extended to incorporate a spatial dimension by using geographical information across provinces to examine whether spatial spillover effects or regional clustering patterns influence the effectiveness of green finance policies and carbon reduction outcomes.